# Multi-unit soft sensing permits few-shot learning in virtual flow metering

### v2

@@ 5.1 Optimization
This objective function is obtained by taking the logarithm of the posterior in (4) and dropping the constant evidence term. These operations do not alter the solution (estimate), but make the objective simpler to optimize numerically. Note that all the terms making up $J$ can be computed analytically for the densities specified in Section 4.
Constrained precision estimation
@@
Mini-batch gradient ascent
@@
$(\hat{\bm{c}},\hat{\bm{t}},\hat{\theta})=\underset{\bm{c},\bm{t},\theta}{\arg\max}~{}J^{r}(\bm{c},\bm{t},\theta\,|\,\mathcal{D}),$ (8)
@@
where $\lambda_{k}$ is the learning rate at step $k$ and the gradient estimate is computed using Algorithm 1. Note that the mini-batch is randomly selected at each iteration. The optimization is stopped when some termination criterion is met, and the final parameter values are returned as the estimate $(\hat{\bm{c}},\hat{\bm{\tau}},\hat{\theta})$, where the precision estimate is retrieved as $\hat{\bm{\tau}}=g(\hat{\bm{t}})$.
@@ 5.2 Context calibration
Let $(\hat{\bm{c}},\hat{\bm{\tau}},\hat{\theta})$ be the parameters estimated from $\mathcal{D}$. We assign the index $M+1$ to a new unit with data $\mathcal{D}_{M+1}$. A simple calibration procedure for $c_{M+1}$ and $\tau_{M+1}$ is to apply the gradient update in (9) with a fixed $\theta=\hat{\theta}$ and $\mathcal{B}=\mathcal{D}_{M+1}$. With this procedure, $K+1$ parameters are calibrated to $|\mathcal{D}_{M+1}|$ data points. Provided that we do not have any prior knowledge about unit $M+1$, it is natural to initialize $c_{M+1}=0$ and $\tau_{M+1}=(\hat{\tau}_{1}+\cdots+\hat{\tau}_{M})/M$.
@@
In the case where $|\mathcal{D}_{M+1}|$ contains very few data points, it may also make sense to fix $\bm{\tau}_{M+1}$ during training to reduce the number of learned parameters.
@@ 5.3 Context parameter analysis
The context parameters in $\bm{c}$ represent latent variables, and cannot be expected to exactly represent physical quantities. However, one may still hope that the learned context parameter space possess certain properties which are commonly held by physical parameter spaces. An important such property is that similar units are represented by similar context parameters. To investigate the validity of beliefs like this, methods for analysing the context parameters are required. In this section we describe an approach which may indicate whether the learned context parameter space is well-behaved; for example in the sense of being smooth.
Laplace approximation
MAP estimation, as presented above, gives a point estimate $(\hat{\bm{c}},\hat{\bm{\tau}},\hat{\theta})$ of the mode of the posterior distribution $p(\bm{c},\bm{\tau},\theta\,|\,\mathcal{D})$. Thus, it does not inform us of the uncertainty in the estimate. In the few-shot learning setting, the marginalized posterior distribution $p(\bm{c}\,|\,\mathcal{D})$ may provide useful information about the context uncertainty given few observations, which may, in turn, be useful for analyzing context parameters. While the exact posterior is hard to compute, a local approximation $q(\bm{c}\,|\,\mathcal{D})$ of $p(\bm{c}\,|\,\mathcal{D})$ can be obtained by Laplace’s approximation [9]:
@@ 6 Empirical study
In the following, a single petroleum well constitutes a unit. The soft sensor that results from applying our proposed model as a VFM is similar to those used in [20, 21] to perform multi-task learning across wells. However, the few-shot learning experiments conducted in this work consider a different learning problem and data availability regime than those presented in said works. Our model architecture also differs by its somewhat less problem-specific architecture as well as its inclusion of a learnable precision parameter which is unique for each unit. Finally, we conduct our study on a larger number of wells than previous studies; to our knowledge, our case study is the largest one in the VFM literature which uses real production data, in terms of the number of wells.
@@ 6.1 Dataset
The dataset consists of data from 80 different petroleum wells, with instrumentation illustrated by Figure 3. The basic system structure and instrumentation are the same for all wells, while physical characteristics (like pipe lengths, diameters, friction coefficients and viscosities) will differ between the wells. Thus, this dataset can be interpreted as an instance of the setting described in Figure 1.
The data take the form of pairs $(x_{ij},y_{ij})$ of observations, where
@@
Here, $u$ denotes choke valve opening, $p^{\textup{WH}}$ and $p^{\textup{DC}}$ denote pressures at the wellhead and downstream the choke valve, $T^{\textup{WH}}$ denotes wellhead temperature, $\eta^{\textup{OIL}}$ and $\eta^{\textup{GAS}}$ denote volumetric oil and gas fractions, $Q^{\textup{GL}}$ denotes volumetric gas lift rate and $Q^{\textup{TOT}}$ denotes total volumetric flow rate. Each data point consists of an average value of each measurement taken over an interval where flow is close to steady-state, detected using the technology described in [40]. Before processing, the observations were shifted and scaled to approximately lie in the unit interval. The measurements of total flow rate include both well tests and MPFM measurements, which our model could address through the use of separate precision parameters for each instrument type as done in [41]. However, for the sake of simplicity, we did not differentiate between the two types of data points in our experiments. The dataset consists 89 417 data points from $M=80$ different wells.
@@ 6.2 Pretraining
The model parameters $(\bm{c},\bm{\tau},\theta)$ can be estimated from data by using the gradient ascent method in (9), or another stochastic gradient-based method, with gradients estimated by Algorithm 1. Since the dimension of $\theta$ exceeds the dimensions of $\bm{c}$ and $\bm{\tau}$ by orders of magnitude for our model, one can expect that most of the complexity of this learning problem can be attributed to the estimation of gradients with respect to $\theta$. Thus, it would be convenient to identify settings in which one can justify freezing $\theta$, instead only estimating $(\bm{c},\bm{\tau})$.
To guide the search for such conditions, one can consider the theoretical convergence rate of the MAP estimator. Assuming that the units are sufficiently similar, and that each unit contributes with a comparable number of data points that does not change with $M$, the theoretical expected learning rate as a function of the number $M$ of units in the dataset will, in a large number of multi-task learning settings, be $\mathcal{O}(1/\sqrt{M})$ [26]. Since this function flattens out as $M$ increases, one should expect diminishing returns from each single new well as their number increases.
We tested this assumption by training and estimating the performance of a model on data from $M=1,5,10,20,40,80$ units. For each value of $M$, the dataset was randomly divided into $\frac{80}{M}$ disjoint groups of $M$ different units. Then, a model was learned from the training data shown in Figure 4 corresponding to each group, using the method described in Section 5. Implementation details are described in Appendix A.
After training, the model was evaluated on a previously unseen test set drawn from the same wells as its training set (as illustrated by Fig. 4). This procedure was repeated 20 times with different random seeds to reduce noise due to well group partitioning, and from stochasticity related to neural network training.
@@
Fig. 5 shows the test set MSE, averaged over all wells and experiment repetitions, as a function of the number of units from which data were available during training. In addition, the figure shows a function in $\mathcal{O}(1/\sqrt{M})$ fitted to this curve using least squares regression. From the figure, one can see that the theoretically predicted convergence rate lies close to the observed one. Furthermore, one can see that the prediction error flattens out significantly between 40 and 80 wells.
Averages over wells for each individual experiment run are shown in Fig. 9 in Appendix C.
@@ 6.3 Few-shot learning performance
Our objective is to demonstrate the use of few-shot learning to achieve high data-driven VFM performance in data-scarce settings, by performing fine-tuning from a model learned from a larger dataset. To this end, the wells were partitioned into two sets: The set of base wells, which provide many data points and are used for model pretraining, and the set of holdout wells, which provide few data points and are used for fine-tuning.
For the holdout wells, the data could not be used as is, since most of the wells shown in Figure 4 are arguably not suffering from data scarcity (in particular, this is true for the wells equipped with an MPFM). To simulate a low-data setting for the holdout wells, we used a method where measurements appear one at a time, sequentially in time, with at least one week in between. For a more detailed description of the dataset generation method, see Appendix B.
In the first step of the experiment, the procedure described in Section 5 was used to learn a base model from a base training set consisting of all available data from 60 randomly chosen base wells. The reason for choosing this particular number of base wells was largely motivated by the need for a holdout set of a size sufficient for robust evaluation of model performance. However, we argue that $M=60$ is situated safely in what can be considered a flat part of Fig. 5, justifying freezing the value $\theta$ learned from the 60 base wells.
Pretraining resulted in a base model with parameters $(\hat{\theta},\hat{\bm{c}},\hat{\bm{\tau}})$. This model was used in the succeeding calibration step, which was performed for each of the 20 holdout wells $k>M$. Here, the neural network parameters $\theta=\hat{\theta}$ were fixed at their value from pretraining, the precision parameter was kept fixed at $\hat{\tau}_{k}=(\hat{\tau}_{1}+\cdots+\hat{\tau}_{M})/M$, while the context parameter $c_{k}$ was calibrated to datasets of increasing size from the holdout well (from 0 up to 10 data points), using the method described in Section 5.2. Further implementation details are given in Appendix A.
@@
Figure 6: Few-shot learning performance. The pretrained models had access to an average of 903 data points from each well.
Fig. 6 describes the development of the mean absolute percentage error (MAPE) on the test set as the number of available data points increases. Here, the blue line shows the median test MAPE across wells, and the colored region shows 50% probability intervals over all wells. The black dashed line indicates the 2.11% MAPE achieved by the base model when evaluated in the same way as the few-shot learning model on the first week of the test set shown in Fig. 4.
@@ 6.4 Analysis of context posteriors
Fig. 7 describes the information gain $D_{\text{KL}}\left({q(\bm{c}\,|\,\mathcal{D}_{k})}\,\lVert\,{p(\bm{c})}\right)$ for holdout datasets $\mathcal{D}_{k}$ of increasing size, as estimated by the procedure described in Section 5.3. The blue line shows the information gain as a function of dataset size, while the shaded blue region shows 50% probability intervals calculated across wells. One can see from the figure that dataset information gain increases steadily with dataset size. Furthermore, the slope of the curve seems to be decreasing, indicating that the effect of additional data points is diminishing.
Figure 7: Information gain of datasets with increasing size.

### v3

@@ 5.1 Optimization
This objective function is obtained by taking the logarithm of the posterior in (4) and dropping the constant evidence term. These operations do not alter the solution (estimate), but make the objective simpler to optimize numerically. Note that all the terms making up $J$ can be computed analytically for the densities specified here.
5.2 Constrained precision estimation
@@ 5.3 Stochastic gradients
5.4 Mini-batch gradient ascent
@@
$(\hat{\bm{c}},\hat{\bm{t}},\hat{\theta})=\underset{\bm{c},\bm{t},\theta}{\arg\max}\leavevmode\nobreak\ J^{r}(\bm{c},\bm{t},\theta\,|\,\mathcal{D}),$ (8)
@@
[…]
@@ 5.5 Context calibration
[…]
@@
In the case where $|\mathcal{D}_{M+1}|$ contains very few data points, it may also make sense to fix $\tau_{M+1}$ during training to reduce the number of learned parameters.
@@ 5.6 Context parameter analysis
The context parameters in $\bm{c}$ represent latent variables, and cannot be expected to exactly represent physical quantities. However, one may still hope that the learned context parameter space possess certain properties which are commonly held by physical parameter spaces. An important such property is that similar units are represented by similar context parameters. To investigate the validity of beliefs like this, methods for analyzing the context parameters are required. In this section we describe two approaches which may indicate whether the learned context parameter space is well-behaved.
5.7 Information gain estimation
[…]
@@ 5.7 Information gain estimation
5.8 Context parameter dimension
For a given unit, the context parameter $c$ can in principle take any value in $\mathbb{R}^{K}$. However, there is no guarantee that all degrees of freedom are used by the model. To better understand how the MTL architecture considered here learns to represent differences between units, one can analyze the intrinsic dimension of $c$.
[…]
To estimate intrinsic dimension in practice, we adapt the method described by [44] to the problem of finetuning $c$. The basis for this method is the definition
$c=c_{0}+Pc^{(k)}$ (11)
where $c_{0}\in\mathbb{R}^{K}$ is a random, fixed vector, $P\in\mathbb{R}^{K\times k}$ a random, fixed, projection matrix, and the vector $c^{(k)}\in\mathbb{R}^{k}$ is learnable and initialized as zero. Dimension constrained learning can then be performed by freezing $\hat{\theta}$ and $\hat{\tau}$ and learning $c$ as defined by (11) with Algorithm 1. Since we have $p(c)=\mathcal{N}(0,I)$, we simplify this procedure by defining $c_{0}=0$.
@@ 6 Empirical study
In the following, a single petroleum well constitutes a unit. The soft sensor that results from applying our proposed model as a VFM is similar to those used in [23, 24] to perform multi-task learning across wells. Another relevant work is [45], which includes an investigation of noise models for neural network-based VFMs. However, the few-shot learning experiments conducted in this work consider a different learning problem and data availability regime than those presented in said works. Our model architecture also differs by its somewhat less problem-specific architecture as well as its inclusion of a learnable precision parameter which is unique for each unit. Finally, we conduct our study on a larger number of wells than previous studies. Studies in the literature typically concern the modeling of a single well [19, 20, 21]. Our model architecture, in contrast, is applied jointly to 80 different petroleum wells. To our knowledge, our case study is the largest one in the VFM literature which uses real production data, in terms of the number of wells (including earlier work on multi-task learning [23]).
6.1 Learning problem
For a single well at a given point in time, it may be possible to derive fairly simple relations between relevant sensor measurements and the total flow. The physics-based models which are used as the basis for VFMs in [19, 20, 21, 24] fall into this category. For the instrumentation shown in Figure 3, such relations often revolve around a model of choked flow which can be derived from Bernoulli’s law:
$Q^{\textup{TOT}}=V(u)C\sqrt{\Delta p}$ (12)
Here, $Q^{\textup{TOT}}$ is the total volumetric well flow, $\Delta p$ is the pressure drop across the choke valve, $V(u)$ is a one dimensional function which depends on the choke valve geometry, and $C$ is a scalar which depends on properties of the fluid mixture which is produced from the well. The pressure drop $\Delta p$ is simple to calculate, and a model of the valve characteristic $V(u)$ may be available from the vendor of the choke valve. If $C$ were easy to estimate, this would then be a fairly simple problem of one dimensional regression. The reason why virtual flow metering is, by no means, a trivial problem, can largely be attributed to the complexity of modeling $C$. Due to the multi-phase nature of the flow which is being modeled, first principles-based modeling approaches typically consist of multiple nonlinear equations which must be jointly solved to provide an estimate of the fluid properties which in the case of (12) is represented by $C$ [46, 47]. The model of $C$ will typically, similarly to the purely data-driven approach which will be used in the following case study, depend on observations of pressure, temperature, flow composition and gas lift rate. The optimal choice of equations for modeling $C$ is, however, strongly dependent on flow conditions, which varies between wells (and which may also vary over time for a given well). Furthermore, wells included in our case study may experience complex flow phenomena like slugging and critical flow, which may not be handled by steady-state first principles models. Thus, first principles-based approaches like those based on (12) are generally not applicable to the multi-unit dataset used here without significant manual work.
Our focus on data-driven models can be further motivated by [24], which compares purely data-driven models based on multi-task learning with a mechanistic modeling approach of a similar (but more complex) structure to (12) in a multi-unit setting. Motivated by the superior performance of the MTL-based data-driven models reported by [24], we do not investigate mechanistic or hybrid approaches in our case study. For further investigation of the advantages and disadvantages of a purely data-driven approach, comparisons with mechanistic models, and for an investigation of how these can be combined into a hybrid VFM, we refer to [24].
@@ 6.2 Dataset
Figure 4: Water cut over time. The gray lines show the development for all 80 wells, and the blue line illustrates the general shape of these curves by showing the development for a single well.
Figure 5: Cumulative distribution of change in water cut between data points, over all 80 wells.
The dataset consists of 89 417 data points distributed across 80 petroleum wells from six oil fields, with instrumentation illustrated by Figure 3. All wells are situated offshore, and produce a mixture of oil, gas and water. The data take the form of pairs $(x_{ij},y_{ij})$ of observations, where
@@
Here, $u$ denotes choke valve opening, $p^{\textup{WH}}$ and $p^{\textup{DC}}$ denote pressures at the wellhead and downstream the choke valve, $T^{\textup{WH}}$ denotes wellhead temperature, $\eta^{\textup{OIL}}$ and $\eta^{\textup{GAS}}$ denote volumetric oil and gas fractions, $Q^{\textup{GL}}$ denotes volumetric gas lift rate and $Q^{\textup{TOT}}$ denotes total volumetric flow rate. Each data point consists of an average value of each measurement taken over an interval where flow is close to steady-state, detected using the technology described in [48]. The observation frequency of these steady-state averaged data points varies depending on instrumentation, but generally appear at frequency on the order of days for MPFM measurements and weeks for well tests.
Out of the 80 wells, 69 employ gas lift and 41 have a multi-phase flow meter installed. Well tests are available for 61 of the wells. The basic system structure and available measurements are the same for all wells, while physical characteristics (like pipe lengths, diameters, friction coefficients and viscosities) differ between the wells. Thus, this dataset constitutes an example of multi-unit data, as described by Figure 1.
Flow conditions also differ across the wells, and many wells are likely to experience complex flow phenomena like slugging or critical flow. Using the rule of thumb which states that critical flow happens when $P^{\textup{DC}}<0.544P^{\textup{WH}}$ [49, 50], we estimate that 67 of the 80 wells experience some critical flow, and 28 of the 80 wells experience critical flow for more than 25% of the available observations. No filtering is done to remove data points subject to such effects, neither in the training nor in the test set.
Before processing, the observations were shifted and scaled to approximately lie in the unit interval. The measurements of total flow rate include both well tests and MPFM measurements, which our model could address through the use of separate precision parameters for each instrument type, as done in [45]. However, for the sake of simplicity, we did not differentiate between the two types of data points in our experiments. We also refer to [45] for a discussion of homoscedastic versus heteroscedastic noise models.
The oil and gas fractions $\eta^{\textup{OIL}}$ and $\eta^{\textup{GAS}}$ are only available from well tests and MPFM measurements. Thus, online observations of these variables will generally not be available to the model when it is used as a VFM. This may cause problems, because fractions tend to change significantly over time. The drift in fractions is apparent from Figure 4, which shows how the water cut (WC), i.e. the ratio between flow rate of water and the total flow rate of all liquids, develops over time for the dataset used in this case study. As indicated by the figure, the water cut typically starts low (often close to zero) and increases steadily during the lifetime of the well, ending up at a high level (often approaching $100\%$). The gas fraction $\eta^{\textup{GAS}}$ also tends to vary over time in a similar, but somewhat less monotonic manner.
However, it is worth noting that the change in fractions over time tends to be slow. This is illustrated by Figure 5, which shows the cumulative distribution (over all wells in the dataset) of the absolute change in water cut between subsequent data points. The blue line shows this distribution for all data, and the red and green lines show the distributions for data which have been down-sampled to have observation frequencies of minimum a week and a month, respectively. For the complete dataset, the approximation error introduced by using the previous available well test seems to be relatively small (for over 60 of the 80 wells, the mean change in WC is below 0.02). The error increases as the observation frequency goes down, but is still fairly low even in the case of monthly observations (where over half of the wells have mean changes below 0.025). In an online implementation of the methods proposed here, one should expect the error introduced by a zeroth-order approximation where one sets the model inputs $\eta^{\textup{OIL}}$ and $\eta^{\textup{GAS}}$ equal to their value at the last available measurement to be relatively small, as long as the observation frequency is not too low. When the observation frequency decreases, however, one should expect model performance to deteriorate accordingly.
@@
P10 P25 P50 P75 P90
CHK median 0.35 0.44 0.58 0.99 1.00
CHK IQR 0.00 0.07 0.15 0.29 0.47
CHK range 0.39 0.51 0.73 0.81 0.86
Pres. drop median 0.01 0.02 0.10 0.25 0.35
Pres. drop IQR 0.01 0.01 0.09 0.17 0.41
Pres. drop range 0.11 0.27 0.44 0.60 0.71
GLR median 0.01 0.04 0.11 0.20 3.18
GLR IQR 0.00 0.01 0.04 0.11 0.91
GLR range 0.04 0.10 0.19 0.40 4.01
Table 1: Summarizing statistics for case study dataset. CHK denotes choke valve opening and GLR denotes gas-to-liquid ratio, both of which are fractions which take values in the unit interval. Pressures are initially given in bar, but are scaled down by a factor of 100 to approximately lie in the unit interval.
Table 1 provides a high-level summary of the dataset by reporting the distribution of and variation in choke valve opening, pressure drop and gas-to-liquid ratio (GLR). The table puts particular emphasis on inter-well variation, and reports the 10th, 25th, 50th, 75th and 90th percentiles across wells of the per-well median, inter-quartile range (IQR) and total range. As can be seen from this table, both operating conditions and variation in operating conditions differ significantly between wells. For instance, the per-well IQR of the choke valve has a lower quartile value (across wells) of 7%, while its upper quartile is situated at 29%. The lower and upper quartiles of the per-well pressure drop range lie at 27 bar and 60 bar, respectively. The per-well median of the gas-to-liquid ratio has lower and upper quartiles at 0,04 and 0,20.
For an in-depth analysis of additional data challenges (including problems relating to well depletion, operating practices and resulting covariate correlation) for a dataset which is largely overlapping with ours, we refer to Chapter 4 of [51].
@@ 6.3 Pretraining
The model parameters $(\bm{c},\bm{\tau},\theta)$ can be estimated from data by using the multi-task learning gradient ascent method in (9), or another stochastic gradient-based method, with gradients estimated by Algorithm 1. Since the dimension of $\theta$ exceeds the dimensions of $\bm{c}$ and $\bm{\tau}$ by orders of magnitude for our model, one can expect that most of the complexity of this learning problem can be attributed to the estimation of gradients with respect to $\theta$. Thus, it would be convenient to identify settings in which one can justify freezing $\theta$, instead only estimating $(\bm{c},\bm{\tau})$.
To guide the search for such conditions, one can consider the theoretical convergence rate of the MAP estimator in the context of MTL. Assuming that the units (which in this case define tasks) are sufficiently similar, and that each unit contributes with a comparable number of data points that does not change with $M$, the theoretical expected learning rate as a function of the number $M$ of units in the dataset will, in a large number of multi-task learning settings, be $\mathcal{O}(1/\sqrt{M})$ [30]. Since this function flattens out as $M$ increases, one should expect diminishing returns from each single new well as their number increases.
We tested this assumption by training and estimating the performance of an MTL model on data from $M=1,5,10,20,40,80$ units. For each value of $M$, the dataset was randomly divided into $80/M$ disjoint groups of $M$ different units, resulting in an MTL dataset with $M$ distinct tasks. Then, a model was learned from the training data shown in Figure 6 corresponding to each group. The neural network was chosen to have large capacity considering the problem complexity; see the appendix for additional implementation details.
After training, the model was evaluated on a previously unseen test set drawn from the same wells as its training set (as illustrated by Figure 6). This procedure was repeated 20 times with different random seeds to reduce noise due to well group partitioning, and from stochasticity related to neural network training.
@@
Figure 7 shows the test set MSE, averaged over all wells and experiment repetitions, as a function of the number of units from which data were available during training. In addition, the figure shows a function in $\mathcal{O}(1/\sqrt{M})$ fitted to this curve using least squares regression. From the figure, one can see that the theoretically predicted convergence rate lies close to the observed one. Furthermore, one can see that the prediction error flattens out significantly between 40 and 80 wells.
Averages over wells for each individual experiment run are shown in Figure 12 in the appendix.
@@ 6.4 Few-shot learning performance
Figure 8: Few-shot learning performance. The x-axis shows the number of available data points for the wells on which few-shot transfer learning is performed. The pretrained models had access to an average of 903 training data points from each well.
In this case study, multi-task pretraining is mainly a means to achieve few-shot transfer learning; our main objective is to demonstrate the use of few-shot learning to achieve high data-driven VFM performance in data-scarce settings, by performing fine-tuning from an MTL model learned from a larger dataset. To this end, the wells were partitioned into two sets: The set of base wells, which provide many data points and are used for MTL model pretraining, and the set of holdout wells, which provide few data points and are used for few-shot transfer learning by context parameter fine-tuning.
For the holdout wells, the data could not be used as is, since most of the wells shown in Figure 6 are arguably not suffering from data scarcity (in particular, this is true for the wells equipped with an MPFM). To simulate a low-data setting for the holdout wells, we used a method where measurements appear one at a time, sequentially in time, with at least one week in between. For a more detailed description of the dataset generation method, see the appendix.
In the first step of the experiment, the multi-task learning procedure described in Algorithm 1 was used to learn a base model from a training set consisting of all available data from 60 randomly chosen base wells. The reason for choosing this particular number of base wells was largely motivated by the need for a holdout set of a size sufficient for robust evaluation of model performance. However, we argue that $M=60$ is situated safely in what can be considered a flat part of Figure 7, justifying freezing the value $\theta$ learned from the 60 base wells.
[…]
Figure 9: Information gain for training datasets of increasing size. The x-axis shows the number of available data points for the wells on which few-shot transfer learning is performed.
@@
Figure 8 describes the development of the mean absolute percentage error (MAPE) on the test set as the number of available data points increases. Here, the blue line shows the median test MAPE across wells, and the colored region shows 50% probability intervals over all wells. The black dashed line indicates the 2.11% MAPE achieved by the base model when evaluated in the same way as the few-shot learning model on the first week of the test set shown in Figure 6.
@@ 6.5 Analysis of context posteriors
Figure 9 describes the information gain $D_{\text{KL}}\left({q(\bm{c}\,|\,\mathcal{D}_{k})}\,\lVert\,{p(\bm{c})}\right)$ for holdout datasets $\mathcal{D}_{k}$ of increasing size, as estimated by the Laplace approximation procedure described above. The blue line shows the information gain as a function of dataset size, while the shaded blue region shows 50% probability intervals calculated across wells. One can see from the figure that dataset information gain increases steadily with dataset size. Furthermore, the slope of the curve seems to be decreasing, indicating that the effect of additional data points is diminishing.
6.6 Context parameter dimension analysis
Figure 10: Scaled test set MSE for different context parameters dimensions. Values are scaled by the MSE achieved for dimension 10 (0.0378), which is comparable to the MSE achieved for $M=80$ in the pretraining experiment (0.0359). Note the logarithmic scaling of the y-axis.
The role of the context parameter dimensionality was analyzed by exchanging the context parameter of the pretrained MTL models with context parameters described by (11), and performing intrinsic dimension estimation as described above. This was done for all 20 pretraining experiment repetitions, with a new projection matrix $P$ being initialized for each well and experiment repetition.
Figure 10 shows the resulting test set MSE, averaged over wells and repetitions, scaled as a fraction of the test set MSE achieved with dimension $k=K=10$, for different group sizes. For all $M>1$, the plot shows that the first few dimensions make the largest contributions to reducing test set error, which flattens out somewhat around dimension 5-6. Still, it is worth noting that the y-axis starts at a fraction which represents rather poor model performance (10 times the best MSE), and that test set error significantly decreases from dimension 6 and all the way up to the full 10 dimensions. Finally, the number of dimensions needed to capture the variation between wells grows with the number of wells which is modeled simultaneously. In the single-unit case $M=1$, the context parameter seems to be of little use.